# Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task

### v1

@@ 6 Results
6.1 Holistic Analysis
Holistic analysis necessitates models with multi-image and multimodal capabilities. Among the models tested, only Gemini 2.0 (flash) (referred to Gemini here) and GPT-4o possess these features.
Table 1 presents the overall accuracy, as well as the accuracy on positive and negative samples separately. As shown, Gemini achieves a slightly higher overall accuracy of 82.2% compared to GPT-4o’s 80.0%. Unlike GPT-4o, Gemini shows a more balanced performance across both positive and negative samples, suggesting a more robust approach to handling cases where no clear rule might exist for the negative examples.
These results suggest that while GPT-4o may excel at recognizing patterns within the positive examples, its lower accuracy on negative samples indicates challenges in distinguishing subtle counter examples. On the other hand, Gemini’s balanced performance reflects greater robustness in addressing the diverse reasoning demands of holistic analysis tasks.
Model Holistic Analysis Deductive Rule Learning Componential Analysis
@@ 6.1 Holistic Analysis
6.2 Deductive Rule Learning
This two-stage approach allows us to evaluate the model’s ability to not only identify patterns but also articulate them as explicit rules, providing insights into its deductive reasoning capabilities. We conducted these experiments using Gemini and GPT-4o. Table 1 summarizes the overall accuracy, as well as the accuracy on positive and negative samples separately.
This time GPT-4o achieves a slightly higher overall accuracy compared to Gemini. Both models demonstrate balanced performance across positive and negative samples. GPT-4o, in particular, excels in classifying positive samples, while Gemini achieves a more consistent performance across both categories.
These results indicate that the two-stage deductive approach enhances the models’ ability to handle the complexities of both positive and negative examples, surpassing the performance seen in the holistic approach. GPT-4o’s superior accuracy on positive samples suggests its effectiveness in identifying and applying explicit patterns, while Gemini’s balanced results highlight its robustness across diverse reasoning scenarios.
@@ 6.3 Componential Analysis
The componential analysis approach provides a framework for evaluating a broader spectrum of vision-language models (VLMs) by decomposing the complex multi-image reasoning task into manageable single-image description tasks. This design facilitates the inclusion of models that lack explicit multi-image processing capabilities but can effectively analyze individual images and process structured textual representations.
We conducted the three-stage componential analysis experiment using six VLMs: Pixtral-12B, Llava-Llama3-8B, Llama-Vision-11B, Llava 7B, GPT-4o, and Gemini. The performance metrics, presented in Table 1, include overall accuracy and accuracy on positive and negative samples.
The results demonstrate a wide range of performance across the evaluated models. GPT-4o and Gemini achieve near-optimal accuracies, with overall accuracies of 92.8% and 93.6%, respectively. Notably, these results surpass human-level performance, which is reported at 91% on this task. This finding highlights the potential of advanced VLMs to excel in tasks traditionally dominated by human reasoning capabilities.
Open-source models, including Pixtral-12B, Llava-Llama3-8B, and Llama-Vision-11B, Llava 7B exhibit significant variability. Llava-Llama3-8B, for example, achieves a high accuracy on negative samples but an exceptionally low accuracy on positive samples, indicating a strong bias toward negative classifications. Similarly, Llama-Vision-11B demonstrates an inverse pattern. These discrepancies suggest that these models struggle to generalize effectively across different classes.
The superior performance of GPT-4o and Gemini indicates that structured, detailed descriptions significantly enhance reasoning capabilities in VLMs. Their ability to surpass human-level performance underscores the importance of integrating multimodal capabilities with robust processing pipelines. Furthermore, performance open-source models highlights the potential for continued advancements in accessible VLMs to achieve substantial progress in complex reasoning tasks.
@@ 6.4 Rule-Based Evaluation
The Rule-Based Evaluation experiment isolates the models’ ability to apply explicitly provided rules to classify query images, bypassing the rule inference step. This approach offers a focused assessment of the rule application capabilities of the models. The results, presented in Table 2, reveal distinct trends across the evaluated models.
Pixtral-12B achieved the highest overall accuracy (83.8%), indicating a strong ability to comprehend and operationalize the provided rules effectively. Its relatively balanced performance on both negative and positive samples highlights its robust application of rules across varying contexts.
@@
The results in Table 2 highlight that while certain models excel in confirming positive matches, their difficulties in handling negative cases expose potential biases in rule application mechanisms. Pixtral-12B’s relatively balanced performance suggests it is more adept at applying rules across diverse scenarios, while models like Llama-Vision-xx and LLaVA-xx may overfit to confirmatory evidence.
By isolating the rule application component, this experiment provides valuable insights into the models’ strengths and limitations. When combined with findings from the Deductive Rule Learning experiments (Section 5.2), this analysis helps to disentangle the factors contributing to model performance, distinguishing between deficits in rule extraction and rule application. Such insights are crucial for guiding improvements in visual-language model reasoning capabilities.
7 Ablation
@@ 7.1 Semantic Similarity Analysis
To further analyze the reasoning process of the best-performing models, GPT-4o and Gemini, we conducted a semantic similarity analysis.This analysis aims to quantify the alignment between the descriptions of the query images and the rules identified for each category. We hypothesized that for correctly classified positive examples, the similarity between the description and rule should be high, while for negative examples, it should be lower.
[…]
7.2 Analysis per Commonsense Category
To better understand the comparative capabilities of GPT-4o and Gemini, we analyzed their performance across ten commonsense reasoning categories in the Bongard Openworld dataset (details in Appendix B.3). This analysis (in Table A.5) highlights the models’ strengths and limitations in handling diverse commonsense reasoning categories. Both models demonstrated robust performance across most categories indicating models’ capability to process high-level conceptual patterns effectively.
Differences in performance were observed in categories requiring more nuanced or contextual understanding, such as "Taste / Nutrition / Food" and "Human-Object Interaction." This suggests that while both models are adept at general reasoning tasks, their architecture and training data influence their performance on specific types of visual information. Notably, Gemini showed strengths in scenarios involving fine-grained object attributes, while GPT-4o performed better in tasks demanding contextual interpretation.
@@ 7.3 Analysis of Incorrectly Classified Samples
In this ablation study, we analyzed the misclassified samples from GPT-4o and Gemini during the Componential Analysis experiments to identify common limitations. The total 15 common test samples were incorrectly identified by both the VLMs. The errors revealed several recurring issues (more details in Table A.7), including challenges with synthetic images that lack natural textures, difficulties in making fine-grained distinctions, and struggles with recognizing specific image types, such as satellite views. Additionally, the models showed limitations in understanding nuanced contextual information, as evidenced by misclassifications related to artistic styles and the broader scene context. These findings highlight that while the models excel in many areas, they still face challenges in handling complex visual details and contextual reasoning, suggesting areas for improvement in VLMs’ robustness and generalization.
@@ 8 Discussion
Our study investigates the application of Vision-Language Models (VLMs) to Bongard Open-World Problems (BPs), providing insights into the computational mechanisms underlying visual reasoning and their alignment with human cognitive strategies. We introduced three evaluation paradigms—Holistic Analysis, Deductive Rule Learning, and Componential Analysis—designed to mirror distinct human approaches to learning and problem-solving. These paradigms allow us to evaluate VLM performance and understand how these models relate to human cognition.
A critical aspect of human scene understanding is the dynamic interaction between the observer and their environment Zelinsky (2013); Malcolm et al. (2016); Gibson (2015). This interactionist view posits that understanding a scene is an active process shaped by behavioral goals, rather than a passive reception of sensory input. This active engagement with the visual world is directly relevant to how both humans and VLMs approach BPs. Rather than processing individual images in isolation, both humans and VLMs benefit from considering the relationships between images in a given set.
Our Holistic Analysis paradigm specifically targets this relational aspect by requiring VLMs to process all images and descriptions simultaneously. This approach allows the model to capture the overall context and directly identify the distinguishing rule, akin to human holistic problem-solving, where the "gist" of a scene is rapidly perceived, guiding subsequent attention and actions Biederman (1987); Li et al. (2002). Such holistic processing provides a crucial initial understanding, which can then be refined through more detailed analysis. This process heavily relies on working memory to maintain the "gist" while further analysis unfolds Baddeley (2012).
Following this initial holistic understanding, human visual information processing is further refined through attentional mechanisms that guide gaze and cognitive processing Posner (1980). While eye movements can reflect attentional shifts, our Deductive Rule Learning paradigm directs the VLMs’ attention on learning the underlying rule prior to evaluating query images. This mirrors human deductive reasoning, where a hypothesis is formed and then tested against new evidence. By focusing on rule inference before rule application, the VLM can store learned rules in long-term memory Squire (1992), allowing for efficient application to unseen instances.
In addition to holistic and deductive strategies, humans often use analytical approaches to break down complex problems. The Componential Analysis paradigm reflects this tendency by decomposing a complex problem into smaller, more manageable components. In this paradigm, VLMs analyze positive and negative examples separately, extracting individual rules for each set before comparing them to identify the distinguishing factor. This process aligns with cognitive theories that distinguish between System 1 (fast, intuitive perception) and System 2 (slow, deliberate reasoning) Kahneman (2011). Object recognition and categorization, a System 1 process involving perceptual mechanisms Marr (2010), precedes the more deliberate System 2 reasoning required to compare and contrast extracted rules. This analytical decomposition facilitates a more precise understanding of the underlying rules.
Our findings, which show that VLMs can surpass human performance on BPs when applying these human-inspired paradigms, underscore the importance of attention and memory mechanisms in visual reasoning. Attention mechanisms, critical for selecting relevant information, are implemented in VLMs through architectures such as those described in Bahdanau et al. (2016) and Vaswani et al. (2017). These mechanisms allow VLMs to selectively focus on relevant image features, similar to human gaze patterns and attentional shifts. Meanwhile, memory mechanisms are essential for encoding visual information, maintaining it during rule formulation, and retrieving it for comparison. This is evident in the VLMs’ ability to store learned rules (long-term memory) and maintain representations of images during analysis (working memory). The interplay between attention and memory is crucial for both human and artificial intelligence, providing deeper insights into the computational foundations of visual reasoning Vaishnav and Serre (2023).
Moreover, our evaluation of both open-source and closed-source models highlights significant performance discrepancies, shedding light on the limitations of current VLMs in handling complex visual reasoning tasks. These limitations underscore the need for further model refinement, particularly in handling multi-image reasoning tasks and improving rule extraction capabilities. Finally, we demonstrate that effective reasoning in complex visual scenarios requires the complementary strengths of VLMs and Large Language Models (LLMs). While VLMs excel at extracting features from images, LLMs play a crucial role in higher-level reasoning tasks, indicating a promising direction for future research where the integration of both types of models can enable more sophisticated problem-solving in visual reasoning tasks.
9 Limitation
Despite the promising results of this study, several limitations must be considered. The Bongard Openworld dataset may not fully capture the complexity and diversity of real-world visual reasoning tasks, potentially limiting generalizability. Additionally, the reliance on pre-extracted rule summaries in some experiments restricts the evaluation of models’ complete reasoning capabilities, particularly in rule extraction. The scalability of large models, such as GPT-4o and Gemini, is also a concern due to their high computational demands, and their ability to generalize to noisy, real-world data remains uncertain. Furthermore, the opacity of model decision-making processes highlights the need for better interpretability, and the risk of overfitting to task-specific knowledge could hinder broader applicability. Lastly, our evaluation primarily focuses on accuracy and semantic similarity, which may not fully capture the nuances of model reasoning. These limitations highlight areas for future research and refinement in applying VLMs to complex reasoning tasks.
@@ 10 Conclusion
This work investigates the effectiveness of Vision-Language Models (VLMs) in solving the challenging Bongard Openworld dataset, a task requiring complex visual reasoning. We proposed three human-inspired paradigms—Holistic Analysis, Deductive Rule Learning, and Componential Analysis—and demonstrated that VLMs, particularly GPT-4o and Gemini, can surpass human-level performance. Our results highlight that VLMs excel when combining detailed image analysis with structured rule-based reasoning. The study also emphasizes the complementary roles of VLMs and Large Language Models (LLMs), with VLMs excelling at visual feature extraction and LLMs crucial for higher-level reasoning. These findings provide valuable insights into model design and evaluation, suggesting a promising direction for future research in complex visual reasoning tasks and real-world applications.
@@ Acknowledgments
Appendix A Broader Relevance
Broader Relevance
The insights gained from this work have broader implications beyond the specific task of solving the Bongard Openworld dataset. As the capabilities of Vision-Language Models (VLMs) continue to evolve, the methodologies and paradigms explored in this paper can contribute significantly to the development of models capable of more advanced reasoning and problem-solving in real-world applications. The ability to tackle complex visual reasoning tasks, such as the Bongard problems, is critical for a range of fields, from autonomous systems and robotics to medical imaging and beyond.
One of the most significant implications of our work is the demonstration that VLMs, when integrated with structured reasoning paradigms, can outperform human-level performance in specific problem-solving scenarios. This opens up new avenues for the deployment of AI systems in environments that require a blend of visual perception and reasoning, such as dynamic real-world settings where an agent must continually adapt to new and diverse scenarios. By surpassing human-level performance in tasks that require complex visual analysis, VLMs could enable more robust and adaptive AI systems that can be deployed in areas ranging from automated decision-making to real-time problem solving.
Furthermore, our work also highlights the critical gap between the reasoning abilities of VLMs vs Large Language Models (LLMs). A combination of both is essential is needed for complex reasoning tasks. While VLMs excel at understanding and processing visual inputs, LLMs bring the reasoning and linguistic abilities necessary to contextualize and apply rules based on these inputs. This integration is essential for advancing research in cognitive modeling, where understanding the interplay between perception, reasoning, and language is central. Our findings encourage further exploration of hybrid architectures that seamlessly blend visual and language models, potentially driving progress in areas such as visual question answering, image captioning, and multimodal learning.
Finally, our detailed evaluation of both open-source and closed-source models contributes to the ongoing dialogue about the accessibility, fairness, and transparency of AI systems. By examining the performance discrepancies across various architectures, we not only illuminate the current limitations of VLMs but also provide actionable insights for future improvements. This work could inspire the development of more efficient and capable models, particularly in the open-source community, and contribute to efforts aimed at making advanced AI more accessible and usable for a wide range of applications.
In summary, this work sets the stage for more advanced, human-like AI systems capable of solving complex visual reasoning tasks, with far-reaching applications in domains such as autonomous systems, healthcare, and cognitive modeling. The methodologies and findings outlined here provide a valuable foundation for future research and development in AI, pushing the boundaries of what is possible with multimodal learning and reasoning.
Appendix B Dataset
B.1 Bongard OpenWorld Dataset

### v2

@@ 6 Results
6.1 Direct Visual Rule Learning
Direct visual rule learning necessitates models with multi-image and multimodal capabilities. Among the models tested, only Gemini 2.0 and GPT-4o possess these features.
Table 1 presents the overall accuracy, as well as the accuracy on positive and negative samples separately. As shown, Gemini 2.0 achieves a slightly higher overall accuracy of 82.2% compared to GPT-4o’s 80.0%. These results suggest that while GPT-4o may excel at recognizing patterns within the positive examples, its lower accuracy on negative samples indicates challenges in distinguishing subtle counter examples. On the other hand, Gemini 2.0’s balanced performance reflects greater robustness in addressing the diverse reasoning demands of direct visual rule learning tasks.
6.2 Deductive Rule Learning
The deductive rule learning paradigm, with its two-stage process of rule extraction and application, offers a more structured approach to visual reasoning compared to the direct visual rule learning. This structured approach allows us to evaluate the model’s ability to not only identify patterns but also articulate them as explicit rules, providing insights into its deductive reasoning capabilities. We conducted these experiments using Gemini 2.0 and GPT-4o. Table 1 summarizes the overall accuracy, as well as the accuracy on positive and negative samples separately.
As shown in Table 1, GPT-4o achieves slightly higher overall accuracy than Gemini 2.0, suggesting greater effectiveness in combining rule extraction and application within this paradigm. While both models demonstrate relatively balanced performance across positive and negative samples, subtle differences emerge. Similar to the previous experiment, GPT-4o exhibits a stronger ability to classify positive samples, and Gemini 2.0 achieves more consistent performance across both categories, suggesting greater robustness across diverse reasoning scenarios.
Comparing these results to those obtained using the direct visual rule learning (Table 1), we can see a clear improvement in overall accuracy for both models. This indicates that the two-stage deductive approach enhances the models’ ability to handle the complexities of both positive and negative test samples, surpassing the performance seen in the holistic approach. This improvement likely stems from the explicit separation of the rule extraction and application processes, allowing the models to focus on each task individually. This suggests that explicitly structuring the reasoning process, even within purely visual tasks, benefits VLMs.
Model Direct Visual Rule Learning Deductive Rule Learning Componential Analysis
@@ 6.3 Componential Analysis
The componential analysis approach evaluates VLMs by decomposing the complex multi-image reasoning task into manageable single-image description tasks. This method separates the perception component (image analysis and description) from the reasoning component (rule derivation and application), relying solely on textual representations for the latter. This isolation of perception and reasoning allows for a deeper understanding of their respective contributions to VLM performance.
Table 1 presents the performance metrics for six VLMs: Pixtral-12B, Llava-Llama3-8B, Llama-Vision-11B, Llava 7B, GPT-4o, and Gemini 2.0. The results reveal significant performance variations, highlighting the crucial role of accurate image descriptions in complex visual reasoning. GPT-4o and Gemini 2.0 achieve near-optimal accuracies (92.8% and 93.6%, respectively), surpassing human-level performance i.e. 91% Wu et al. (2024). This demonstrates the potential of advanced VLMs when provided with high-quality, structured descriptions. Their strong performance suggests that these models excel at leveraging textual information for rule derivation and application—the reasoning component.
In contrast, the open-source models exhibit greater performance variability. While Pixtral-12B and Llava 7B achieve moderate, balanced performance, Llava-Llama3-8B and Llama-Vision-11B show a striking bias: one performs well on negative samples but poorly on positive samples, while the other exhibits the opposite trend. This suggests a significant challenge in generalization for these models, likely stemming from inconsistencies or inaccuracies in their generated descriptions—the perception component.
The stark difference in performance between the closed- and the open-source models underscores the importance of robust image description generation. The near- or super-human performance of GPT-4o and Gemini 2.0 highlights the potential of VLMs when both perception and reasoning components are strong. The challenges faced by the open-source models, suggest that future research should prioritize improving the perception modules of these models to achieve human-like reasoning. Further investigation is needed to understand the specific factors contributing to the open-source models’ difficulties in generating consistent and accurate descriptions.
7 Ablation
@@ 7.1 Rule-Based Evaluation
The Deductive Rule Learning paradigm (Section 6.2) combines rule extraction and application. To isolate the latter, we designed an ablation study focusing solely on rule application, thus disentangling perception from reasoning. This allows us to directly assess the models’ reasoning capabilities independent of their perception modules.
Specifically, we used the rule summaries generated by GPT-4o during the rule extraction phase. For each Bongard problem, we provided the open-source models with two inputs: 1) the corresponding rule summary, and 2) the query image. The models were instructed to classify the query image based on the provided rule extracted using GPT-4o. This isolates the models’ ability to understand and apply a given rule, independent of their capacity to extract it.
Table 2 presents the results. Pixtral-12B achieved the highest overall accuracy (83.8%), demonstrating strong rule comprehension and application. Its balanced performance suggests robust rule application across contexts, contrasting with some other models exhibiting a bias towards confirming positive matches. While some achieve high positive accuracy, they struggle with negative cases, suggesting difficulty identifying when a rule does not apply. Pixtral-12B’s balanced performance suggests greater adeptness, while models like Llama-Vision-xx and LLaVA-xx may over-rely on confirmatory evidence.
Comparing the performance of Pixtral-12B and the Llava models with their componential analysis results reveals a crucial insight. Pixtral-12B maintains relatively consistent performance across both evaluations. However, the Llava models’ accuracy drops substantially in the componential analysis compared to the rule-based evaluation. This performance gap strongly indicates that the Llava models’ primary bottleneck lies in the perception stage – the ability to generate accurate and comprehensive image descriptions – rather than in their reasoning capabilities (rule application), which are more effectively assessed when pre-extracted rules are provided.
By isolating rule application, this ablation provides insights into the open-source models’ strengths and limitations in reasoning. Combined with the Deductive Rule Learning results, this analysis disentangles factors contributing to overall performance, distinguishing between deficits in rule extraction (a proxy for perception) and application (pure reasoning).
@@
7.2 Analysis per Commonsense Category
To better understand the comparative capabilities of GPT-4o and Gemini 2.0, we analyzed their performance across ten commonsense reasoning categories – provided in the Bongard Openworld dataset (details in Appendix A.3.3). This analysis highlights the models’ strengths and limitations in handling diverse commonsense reasoning categories (refer to Table A.3). Both models demonstrated robust performance across most categories indicating models’ capability to process high-level conceptual patterns effectively.
Differences in performance were observed in categories requiring more nuanced or contextual understanding, such as “Taste / Nutrition / Food” and “Human-Object Interaction.” This suggests that while both models are adept at general reasoning tasks, their architecture and training data influence their performance on specific types of visual information. Notably, Gemini 2.0 showed strengths in scenarios involving fine-grained object attributes, while GPT-4o performed better in tasks demanding contextual interpretation.
@@ 7.3 Semantic Similarity Analysis
To further analyze the reasoning process of the best-performing models, GPT-4o and Gemini 2.0, we conducted a semantic similarity analysis. This analysis aims to quantify the alignment between the perception (image description) and the reasoning component (identified rule). We hypothesized that for correctly classified positive examples, the similarity between the description and rule should be high, reflecting a strong connection between perception and the derived rule. Conversely, for negative examples, the similarity should be lower, as the description should not align with the rule.
Using cosine similarity of sentence embeddings generated using OpenAI embeddings, we computed the mean and standard deviation of the semantic similarity scores for both models and categories. Both models exhibit higher similarity (see Table A.4) for positive examples compared to negative ones, confirming that correct classifications align with the identified rule. However, the small difference in similarity ($\sim$ 6%) suggests that the negative samples were designed to be closer to positive samples to make them challenging. This highlights the subtle perceptual distinctions required for correct classification in these challenging cases.
@@ 7.4 Analysis of Incorrectly Classified Samples
Here we analyzed the misclassified samples from GPT-4o and Gemini 2.0 during the Componential Analysis experiments to identify common limitations. The total 15 common test samples were incorrectly identified by both the VLMs. The errors revealed several recurring issues (more details in Table LABEL:tab:tasks). We found eleven cases of errors in rule extraction; more specifically, four cases of being too general, one case of a missed property, five cases of the missing required object, two cases of focusing on a wrong property. We also found three cases of weak similarity-based reasoning disregarding some parts of the rule. These findings highlight that while the models excel in many areas, they still face challenges in handling complex visual details and reasoning, suggesting areas for improvement in VLMs’ robustness and generalization.
@@ 8 Discussion
This study investigates how VLMs tackle Bongard Problems (BPs), offering insights into their visual reasoning mechanisms and alignment with human cognitive strategies. We introduced three evaluation paradigms, each mirroring a distinct human problem-solving approach. These paradigms allow us to evaluate VLM performance and dissect the contributions of different cognitive processes to their visual reasoning abilities.
Human scene understanding is an active, interactive process Zelinsky (2013); Malcolm et al. (2016); Gibson (2015). This interactionist perspective is key to understanding how both humans and VLMs approach BPs. Rather than passively receiving sensory input, both benefit from actively considering the relationships between images. Our Direct Visual Rule Learning paradigm directly probes this, requiring simultaneous processing of all images, mirroring holistic human problem-solving where a rapid “gist” perception guides further analysis Biederman (1987); Li et al. (2002). This initial holistic understanding is then typically refined through focused attention Posner (1980). Our Deductive Rule Learning paradigm simulates this by prompting VLMs to first learn the underlying rule before evaluating query images (more in Appendix A.2). This mimics human deductive reasoning where a hypothesis is formed and then tested.
Humans also frequently employ analytical strategies, breaking down complex problems. Our Componential Analysis paradigm reflects this, with VLMs analyzing positive and negative examples separately before comparison. This aligns with dual-process theories of cognition, where intuitive System 1 - perception, precedes deliberate System 2 - reasoning Kahneman (2011); Marr (2010). Critically, this paradigm allows us to isolate the perception component (image description) from the reasoning component (rule derivation and application), revealing the distinct roles these play in VLM performance.
Importantly, our three paradigms can be seen as progressively more detailed instantiations of a Chain-of-Thought (CoT) approach (more details in Appendix A.6.3). Direct Visual Rule Learning implicitly requires holistic image relationship consideration, Deductive Rule Learning adds explicit rule extraction and Componential Analysis represents the most granular CoT, decomposing images into separate descriptions. This progressive increase in detail is mirrored by increased performance, suggesting that structuring visual reasoning benefits VLMs. Unlike typical CoT prompting, our approach uses structured visual and textual representations in multiple stages, offering a way to induce step-by-step reasoning.
Our results show that VLMs can surpass human performance on BPs given high-quality, structured descriptions, notably in the Componential Analysis paradigm. This suggests that mitigating perceptual challenges allows VLMs to effectively use textual representations for sophisticated reasoning. However, the substantial performance gap between open- and closed-source models reveals critical limitations, especially in the perception component. The stark contrast between rule-based evaluation (bypassing perception) and Componential Analysis (requiring it) underscores this bottleneck. For instance, Llava models exhibited a significant performance drop when generating descriptions compared to using pre-extracted rules.
9 Limitations
While this work offers valuable insights, it’s crucial to acknowledge its limitations. Our evaluation relies on the Bongard OpenWorld dataset, which may not fully capture the diversity of real-world visual reasoning scenarios. Exploring more diverse datasets is essential for assessing generalizability. The Componential Analysis paradigm, while insightful, is inherently tied to the quality of generated image descriptions, which are often imperfect, especially for open-source VLMs. Furthermore, the computational demands of state-of-the-art models raise concerns about scalability and accessibility. Finally, our quantitative metrics may not fully capture the nuances of model reasoning.
Future work should prioritize improving open-source VLMs’ perception modules, exploring more diverse datasets, and developing richer evaluation metrics to provide deeper insights into model reasoning. Integrating VLMs and LLMs to leverage their complementary strengths is a promising direction for achieving more human-like reasoning.
@@ 10 Conclusion
This work presents a cognitively-inspired framework for evaluating the visual reasoning capabilities of Vision-Language Models (VLMs) on natural image-based Bongard Problems. By systematically dissecting the interplay between perception and reasoning, we introduce three structured paradigms—Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—that progressively enforce step-wise reasoning. Our findings reveal that advanced, closed-source VLMs (GPT-4o, Gemini 2.0) achieve near-superhuman performance, particularly when provided with high-quality image descriptions, whereas open-source models exhibit a significant performance gap due to deficiencies in their perception modules. The ablation study further highlights that perception, rather than reasoning, is the primary bottleneck, as open-source models can apply extracted rules effectively when given accurate descriptions. These results underscore the critical role of robust multimodal perception in advancing generalizable and interpretable visual reasoning models. Addressing this challenge requires enhancing visual representations, improving multimodal alignment, and designing architectures that enforce structured, step-wise reasoning akin to Chain-of-Thought prompting. Additionally, integrating cognitive principles from human problem-solving strategies could provide a pathway toward developing more robust, transparent, and human-aligned visual reasoning systems.
@@ Acknowledgments
Appendix A.1 Broader Relevance
This study offers insights with broader implications for the development of more robust and human-like AI systems. Our cognitive inspired evaluation paradigms — Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—provide valuable tools for assessing and understanding the strengths and limitations of Vision-Language Models (VLMs) on complex visual reasoning tasks. The insights gained from applying these paradigms to the Bongard Open-World Dataset extend beyond this specific task, contributing to the development of VLMs capable of advanced reasoning in real-world applications.
Our key finding—that the perception module (image description) is a critical bottleneck for open-source VLMs—has significant implications for future research. Improving the ability of these models to generate accurate and comprehensive image descriptions is crucial for bridging the performance gap with closed-source models and unlocking their full potential for complex visual reasoning.
The demonstration that advanced VLMs (GPT-4o and Gemini 2.0) can achieve near- or even super-human performance on BPs, particularly when provided with high-quality descriptions, underscores the potential of these models for tasks requiring sophisticated visual understanding. This reinforces the importance of continued research into multimodal architectures that effectively integrate robust perception with advanced reasoning capabilities.
Finally, our comparative evaluation of open- and closed-source models contributes to the broader discussion about AI accessibility and transparency. By highlighting the performance disparities, we identify specific areas for improvement in open-source models, promoting the development of more accessible and capable AI systems. This work provides a foundation for future research aimed at bridging the gap between human and machine visual reasoning, paving the way for more robust and reliable AI in diverse real-world applications.
Appendix A.2 Attention and Memory in Visual Reasoning
While our study primarily focuses on the interplay between perception and reasoning, the roles of attention and memory are also implicitly present in our paradigms. The Direct Visual Rule Learning paradigm, by requiring simultaneous processing of all images, likely engages VLMs’ “visual attention” mechanisms to identify salient features and relationships across the image set. This aligns with research on human attention, which suggests that holistic scene processing relies on rapid perception and attentional guidance Biederman (1987); Li et al. (2002). These mechanisms allow VLMs to selectively focus on relevant image features, mirroring human gaze patterns and attentional shifts. Similarly, the Deductive Rule Learning paradigm relies on the model’s ability to “memorize” the extracted rule for later application to query images, a process that likely involves both working memory Baddeley (2012) and long-term memory storage Squire (1992). Memory mechanisms are essential for encoding visual information, maintaining it during rule formulation, and retrieving it for comparison. This is reflected in the VLMs’ ability to store learned rules (long-term memory) and maintain representations of the images during analysis (working memory). Although we did not directly manipulate or measure these cognitive processes, their involvement is undeniable, and future research could explore these aspects more explicitly, perhaps by analyzing attention maps Bahdanau et al. (2016); Vaswani et al. (2017) or investigating the memory representations learned by VLMs during rule extraction. The interplay between attention and memory, crucial for both human and artificial intelligence, provides valuable insights into the computational underpinnings of visual reasoning Vaishnav and Serre (2023).
Appendix A.3 Dataset
A.3.1 Bongard OpenWorld Dataset